# How Far are VLMs from Visual Spatial Intelligence? A Benchmark-Driven Perspective

### v1

@@ 3.4 Training Strategy Optimization
VSR goes beyond perceiving static structures, typically requiring dynamic, multi-step reasoning grounded in commonsense knowledge. This multi-step nature makes it a natural fit for reinforcement learning, especially Group Relative Policy Optimization (GRPO) [159], which has already shown success in enhancing the textual reasoning and generation capabilities of LLMs and VLMs. Inspired by this success, a growing body of work [160, 161, 162, 30, 163, 164, 28] is now investigating RL’s potential for spatial reasoning tasks. However, unlike mathematical reasoning [24] or code generation [25] tasks, where correctness can be directly verified, spatial reasoning lacks a clear, well-defined reward signal. This makes it difficult for RL to acquire spatial reasoning abilities effectively, leaving the application of RL to VSR a challenging and open problem.
To deal with such challenges, current works generally adopt two strategies. The first category [163, 28, 162, 164] adopts a two-stage paradigm, in which a Supervised Fine-Tuning (SFT) stage is introduced before Reinforcement Learning Fine-Tuning (RLFT) to provide explicit supervision for spatial reasoning. The SFT stage serves as a warm start for training, while the RLFT stage empowers models with strong generalization ability. For example, RoboRefer [163] first trains VLMs with depth, spatial understanding data, and instruction tuning at the SFT stage and further fine-tunes VLMs with multi-step reasoning data at the RLFT stage. The second category adopts a one-stage paradigm, which focuses on designing task-specific reward functions tailored for spatial reasoning in RL training. Considering the invariance and variations between original and flipped images, Wang et al[161] extend GRPO to Spatial-GRPO, comparing rewards between the original and flipped groups and penalizing the group with a higher score. SpaceR [30] explicitly constructs an object-centric map and defines a reward function based on this map to establish quantitative feedback for spatial understanding. In the video spatial reasoning domain, Li et al[160] incorporate additional explicit IoU-based and recall-based rewards into GRPO. By training with different combinations of reward functions in GRPO, these methods achieve improved performance and generalization ability compared with SFT-trained variants.
@@ 3.5 Inference Strategy Enhancements
Internal enhancement. Internal enhancement refers to leveraging pretrained world knowledge inherent in LLMs without injecting additional external knowledge. According to the objects of enhancement, we divide this into three categories. (1) Multimodal Chain-of-Thought (CoT). Prompting LLMs to conduct several steps of linguistic reasoning has shown effectiveness [165, 166]. Inspired by this, multimodal CoT has been introduced for VSR tasks, enabling joint reasoning over visual and textual spaces and thereby enhancing spatial awareness during intermediate reasoning. To elicit and enhance the reasoning steps, previous works [167, 103] propose to visualize the intermediate steps of reasoning. For instance, Wu et al[167] prompts the pretrained VLMs to generate reasoning traces $z_{1....i}$ and spatial visualizations $v_{1....i}$ in a separate manner as follows:
@@
Inspired by ReAct [168], Yao et al[103] propose to generate intermediate spatial visualizations and reasoning traces simultaneously as follows:
@@
Both methods aim to make reasoning steps interpretable. To deal with the lack of intermediate supervision for spatial CoT, Visual CoT [169] collects a dataset consisting of detailed spatial reasoning steps and bounding boxes of Regions of Interest (RoI), serving as ground-truth for CoT supervision. Based on the dataset, the VLMs can be trained to extract visual tokens from the RoIs and enhance reasoning steps based on both RoIs and the entire input, following the ground-truth chain of thought. This design enables the model to dynamically focus on the relevant regions during spatial reasoning steps. (2) Scene Representation. Inspired by the hippocampus’s central role in contextual memory [170], some methods [31, 171] propose to generate internal representations of spaces and conduct reasoning based on these. Given a video input and a question, Yang et al[31] first construct a cognitive map consisting of object center positions in a grid format. Then the predicted map is leveraged to answer the question, achieving robust results regarding local distance awareness. Similarly, when exploring scenes in embodied tasks, a powerful internal representation is also useful to maintain past and current observations. For example, some works [172, 173] adopt 3D voxels as internal scene representation, on which the occupancy and exploration status are updated first. Afterwards, voxels are projected onto a 2D semantic map [172]. For the unexplored areas, this map is utilized to predict probabilities over three candidate directions and to estimate whether further exploration is worthwhile based on current observations. (3) Attention re-distribution. During reasoning, there exists an imbalance between visual and textual attention. Specifically, image tokens constitute 90% of the input sequence but only account for 10% of the total attention, leading to misalignment between the actual spatial layout of objects in images and instruction prior in texts [124]. To deal with this imbalance, AdaptVis [124] introduces a dynamic attention redistribution strategy that leverages output logits as guidance. When the guidance is low, the attention distribution is smoothed, encouraging the model to explore a broader range of inputs. In contrast, when the guidance signal is high, the attention distribution is sharpened, directing the model to focus on key objects.
@@
External enhancement. External enhancement supplements VLMs with knowledge beyond their pretrained parameters to understand illogical or uncommon spatial relationships. One representative direction is the use of multi-agent system. Marsili et al. [174] designs a cooperative system where different agents interact in a code policy manner: API generation agents first decompose the query into sub-questions, while program synthesis agents receive these sub-questions and generate code to solve each problem. Another line of work is Retrieval Augmented Generation (RAG). For example, Yu et al[175] retrieves templates for subjects-objects relationships as well as other spatial descriptions, and then integrates these templates into the inference stage via in-context learning.
@@ 4.1.1 Attributes
Attributes of objects refer to characteristics that remain relatively stable and are unlikely to change in the short term. These typically include shape, color, size, and quantity. The shape of an object is essential for distinguishing between different objects and understanding scenes. Object shape recognition tasks typically begin with simple 2D geometric figures, which include two types of task setups: whole-shape recognition [176, 177] and sub-shape recognition [178]. MM-Bench [179] increases the difficulty by requiring models to recognize object shapes in real images. Additionally, some datasets use multi-view images as input, such as MMSI-Bench [180], which asks the model to deduce an object’s shape from multiple views, synthesizing the information to infer the shape from a fixed viewpoint. Color also plays a critical role in object recognition, and color-related tasks are generally simpler, requiring the model to identify the color of a specific object [179, 181, 182]. In contrast, perceiving the size of an object is more complex. Some studies present two objects for comparison, asking the model to determine which is taller, shorter, wider, narrower, larger, or smaller. Other research requires the model to estimate an object’s size quantitatively, which involves using commonsense knowledge for direct measurement [31, 27, 183]. In OMNI3D-Bench [184], a reference object’s physical measurement is provided as a standard. Object quantity recognition [185, 176, 181, 186, 187, 188, 189, 34] is one of the more challenging tasks in static attribute understanding. It depends on distinguishing between objects based on attributes such as shape, color, and size. This makes hallucination phenomena more pronounced in tasks involving object quantity recognition.
@@ 4.1.2 States
In contrast to attributes, states refer to characteristics of objects that are subject to change. Common state attributes include an object’s posture, orientation, and its open/closed status. In OpenEQA [182], the task requires the model to estimate the open/closed state of a microwave door, while other studies focus on determining an object’s front-facing orientation in space [32, 181, 190]. Additionally, some research [191, 188] introduces a temporal dimension, requiring the model to understand how states evolve over time. For instance, this might involve describing changes in an object’s rotational angle from the past to the present, or transformations in its shape resulting from specific actions [192]. While the perception of state attributes typically involves describing the properties of a single object or category of objects, and thus falls under basic perception, it is more complex than understanding static attributes. This complexity arises from the diversity of state attributes and the need for commonsense reasoning.
@@ 4.2.1 Static Understanding
Spatial Relationships. Early works in the paper VSR[193] identified 66 types of spatial relationships e.g., under, in front of, facing. The test data consists of image-text pairs, where the text includes a description and a label indicating correctness, such as: <Caption: The cow is ahead of the person. Label: False. >. This requires VLMs to first localize the two objects in the image and then determine whether the abstract relation ‘ahead of’ holds. Later studies introduced additional positional relationships, such as occlusion, reach, and containment. The difficulty in these tasks often lies not in object recognition but in understanding specific spatial relationships. The model must establish connections between the spatial structural relations between objects and the corresponding relation terms.
@@
Object Localization. The object localization task requires the model to accurately identify the position of a specified target within an image. The specific localization methods include detecting the precise spatial coordinates of the target object, determining which bounding box best fits the target, or estimating the coordinates of key points on the target. The BLINK [194] test sample provides an image along with a labeled bounding box for the target object, where the model chooses the bounding box that best aligns with the target from multiple options. The SITE [187] involves the model locating a specific object based on a description or hint and estimating the coordinates of its 2D bounding box. Furthermore, the STI-Bench [195] requires the model to localize the 3D bounding box of the target object within a video. Object localization tasks are also included in other benchmarks [182, 179, 34, 29, 182]. The challenge in these tasks lies in the model’s need to possess strong object detection capabilities while also handling various complexities in the background and interference, such as partial occlusion and similar objects.
@@
Distance Measuring. Distance measuring tasks aim to estimate the distance between objects in an image or the distance from an object to the camera, which is also referred to as depth estimation. Distance measuring typically involves two forms: qualitative judgment of proximity and quantitative numerical estimation. For instance, in the paper BLINK [194], the task requires the model to determine which of two specified points is closer to the camera; in SPAR-Bench [196], the task asks the model to judge which object is closer to or farther from a reference point (such as another object or an observer) based on the 3D center point position of the objects. In addition to qualitative distance analysis, SpatialVLM [26] employs an automated data construction pipeline to obtain distance data, enabling quantitative distance estimation. In the paper SpatialBench [34], desktop-level distance estimation is introduced for use in embodied scenarios. Subsequently, many studies have incorporated distance measurement as a crucial task in spatial reasoning [197, 27, 198, 199, 154, 195, 31, 200, 186, 201, 184, 183, 202, 203, 185, 187]. The challenge of this task lies in the need for the model to infer spatial relationships using implicit cues such as structural information in the scene, object size ratios, and physical knowledge. Furthermore, varying camera parameters may introduce misguidance in distance perception.
@@
Spatial Compatibility Reasoning. Spatial compatibility reasoning tasks require VLMs to determine whether one object can fit the spatial properties or another after thoroughly understanding the object’s shape, size, and spatial relationships. In real-world scenarios, this includes determining whether an object can fit into a given space [182] (’Can another cookie jar fit on the cookie jar shelf?’). This task assesses the model’s reasoning abilities concerning spatial constraints and physical adaptation relationships. Additionally, OmniSpatial [33] introduces reasoning about motion processes, such as determining whether an object can pass through a narrow passage. Some synthetic datasets also require the model to understand layout constraints and graphical combination rules in a 2D space, enabling it to provide fitting solutions. The difficulty of this task lies in the model’s need to not only perceive the size and shape of objects but also reason about physical world common sense, such as “an object too large cannot fit into a small space” or “some uniquely shaped objects cannot be stacked together”.
@@
Trajectory Description. The goal of the trajectory description task is to identify and describe the dynamic changes in the object or camera pose over a time sequence based on visual inputs, generating structured language descriptions. In this context, the trajectory refers to the continuous recording of the camera’s pose over time. The descriptions can be divided into two categories based on the observer (camera viewpoint) and the observed world (objective world): one describes the camera’s orientation and pose, and the other describes the objects within the observed world. In STI-Bench [195], the model is required to infer the position and orientation of the camera at a specific time based on video input and initial pose information. In the Ego-ST Bench [204], the test requires the model to qualitatively describe changes in direction, such as “turn left first and then turn right”. Additionally, multi-view reasoning tasks [180, 33, 203, 194, 195, 154, 205, 198, 206] involve numerous estimations of quantified camera parameters. Another category of trajectory description pertains to the observed world, involving tasks like recognizing changes in the direction of other objects in the video, such as “the car changes from moving straight to turning left”. In the VSI-Bench [31], a task is designed that requires the model to determine the order of appearance of objects in a spatiotemporal sequence. Similarly, in the SPACE [199], multiple landmarks are introduced, and the model is required to determine their sequence. The challenges in these tasks arise from two key issues: first, the model must infer 3D spatial from 2D RGB inputs, which involves geometric principles such as perspective projection, a problem that cannot be resolved by simple image feature matching alone. Second, the estimation errors accumulate as the video sequence length increases, requiring the model to have dynamic error correction capabilities. This is a significant challenge in video pose estimation.
@@
Velocity and Acceleration Estimation. In dynamic spatial reasoning tasks, models are required to estimate motion parameters, such as velocity and acceleration, which are crucial for subsequent planning and decision-making. The OmniSpatial [33] requires models to qualitatively assess the speed and acceleration of objects, like “Which car, A or B, is moving faster?” or “Is the acceleration of the ball on the slope increasing?”. In contrast, STI-Bench directly asks models to estimate the numerical values of velocity and acceleration from video, such as “What is the average speed of the camera?” or “How quickly is the ball accelerating?”. Additionally, STI-Bench [195] categorizes tests into three different environments: desktop, indoor, and outdoor, based on varying accuracy requirements for motion parameter estimations. In DynSuperCLEVR [207], the test introduces two additional settings: future prediction and counterfactual reasoning. Future prediction requires the model to infer the future motion state of an object based on its current speed and acceleration, such as “How fast will the blue sphere be moving after 2 seconds?”. Counterfactual reasoning, on the other hand, asks the model to reason about hypothetical scenarios where speed or acceleration changes, such as “If the green cylinder’s acceleration were doubled, when would it collide with the wall? The task of estimating velocity and acceleration requires the model to integrate displacement changes and temporal information within a 3D space while relying on an understanding of physical laws. The challenges of this task arise from two key issues: first, velocity and acceleration are multi-stage dynamic attributes. Estimating speed involves calculating displacement, followed by the estimation of acceleration, which creates a longer inference chain. Second, counterfactual reasoning and future prediction tasks require the model to reconstruct scenarios, involving complex “what-if” logic chains that are more difficult than simple factual queries.
@@
Spatial Imagination. In this paper, we define spatial imagination tasks as those require reasoning from a hypothetical viewpoint different from the input perspective, given a visual input. It is important to note that some works [33, 183, 208] refer to the camera viewpoint as the “egocentric perspective”, while others [186, 209] refer to it as the “allocentric perspective.” To avoid ambiguity, we use the term “hypothetic perspective” to distinguish it from the camera viewpoint, which is the perspective directly provided to the model. The inputs can be a single image, requiring the model to reason from a hypothetical perspective. For example, “If the waiter is on the side of the vase, then who is on his left?” [33] The SQA3D requires the model to first locate a hypothetical viewpoint in the video based on the description and then perform a question-answering task, such as “Sitting at the edge of the bed and facing the sofa, can I go straight to the coffee table in front of me?” The MINDCUBE [171] considers reasoning from limited views, aiming to assess the model’s ability to mentally model space, such as “If you move from view 1 to view 2, what is the furthest from you?” There is a significant body of research [187, 31, 32, 196] in this area, some of which also accounts for spatial understanding differences across various viewpoints and language conventions [209], or conducts reasoning from synthetic collections of images [210]. The naming of spatial imagination tasks varies across studies, SQA3D [211] labeling it as situational question-answering, and SRBench [210] referring to it as mental rotation. Spatial imagination tasks are not simply direct perception of static object attributes but require the model to construct spatial representations of objects in the mind and perform viewpoint transformations, involving dynamic operations on spatial structures. Thus, the challenge of spatial imagination lies in the need of the model to overcome the limitations of the input perspective, build accurate topological and spatial structural relationships from limited viewpoint inputs and imagine visual cues from the hypothetical perspective.

### v2

@@ 3.4 Training Strategy Optimization
VSR goes beyond perceiving static structures, typically requiring dynamic, multi-step reasoning grounded in commonsense knowledge. This multi-step nature makes it a natural fit for reinforcement learning, especially Group Relative Policy Optimization (GRPO) [guo2025deepseek], which has already shown success in enhancing the textual reasoning and generation capabilities of LLMs and VLMs. Inspired by this success, a growing body of work [li2025videochat, wang2025svqa, zhao2025embodied, ouyang2025spacer] is now investigating RL’s potential for spatial reasoning tasks. However, unlike mathematical reasoning [albalak2025bigmathlargescalehighqualitymath] or code generation [pennino2025reasoning] tasks, where correctness can be directly verified, spatial reasoning lacks a clear, well-defined reward signal. This makes it difficult for RL to acquire spatial reasoning abilities effectively, leaving the application of RL to VSR a challenging and open problem.
To deal with such challenges, current works generally adopt two strategies. The first category [su2025pixel, zhou2025roborefer, zhao2025embodied, ma2025spatialreasoner] adopts a two-stage paradigm, in which a Supervised Fine-Tuning (SFT) stage is introduced before Reinforcement Learning Fine-Tuning (RLFT) to provide explicit supervision for spatial reasoning. The SFT stage serves as a warm start for training, while the RLFT stage empowers models with strong generalization ability. For example, RoboRefer [su2025pixel] first trains VLMs with depth, spatial understanding data, and instruction tuning at the SFT stage and further fine-tunes VLMs with multi-step reasoning data at the RLFT stage. The second category adopts a one-stage paradigm, which focuses on designing task-specific reward functions tailored for spatial reasoning in RL training. Considering the invariance and variations between original and flipped images, Wang et al[wang2025svqa] extend GRPO to Spatial-GRPO, comparing rewards between the original and flipped groups and penalizing the group with a higher score. SpaceR [ouyang2025spacer] explicitly constructs an object-centric map and defines a reward function based on this map to establish quantitative feedback for spatial understanding. In the video spatial reasoning domain, Li et al[li2025videochat] incorporate additional explicit IoU-based and recall-based rewards into GRPO. By training with different combinations of reward functions in GRPO, these methods achieve improved performance and generalization ability compared with SFT-trained variants.
@@ 3.5 Inference Strategy Enhancements
Internal enhancement. Internal enhancement refers to leveraging pretrained world knowledge inherent in LLMs without injecting additional external knowledge. According to the objects of enhancement, we divide this into three categories. (1) Multimodal Chain-of-Thought (CoT). Prompting LLMs to conduct several steps of linguistic reasoning has shown effectiveness [huang2022language, suris2023vipergpt]. Inspired by this, multimodal CoT has been introduced for VSR tasks, enabling joint reasoning over visual and textual spaces and thereby enhancing spatial awareness during intermediate reasoning. To elicit and enhance the reasoning steps, previous works [wu2024mind, li2025imagine] propose to visualize the intermediate steps of reasoning. For instance, Wu et al[wu2024mind] prompts the pretrained VLMs to generate reasoning traces $z_{1....i}$ and spatial visualizations $v_{1....i}$ in a separate manner as follows:
@@
Inspired by ReAct [yao2023react], Yao et al[li2025imagine] propose to generate intermediate spatial visualizations and reasoning traces simultaneously as follows:
@@
Both methods aim to make reasoning steps interpretable. To deal with the lack of intermediate supervision for spatial CoT, Visual CoT [shao2024visual] collects a dataset consisting of detailed spatial reasoning steps and bounding boxes of Regions of Interest (RoI), serving as ground-truth for CoT supervision. Based on the dataset, the VLMs can be trained to extract visual tokens from the RoIs and enhance reasoning steps based on both RoIs and the entire input, following the ground-truth chain of thought. This design enables the model to dynamically focus on the relevant regions during spatial reasoning steps. (2) Scene Representation. Inspired by the hippocampus’s central role in contextual memory [nadel2008hippocampus], some methods [VSI-Bench, MINDCUBE] propose to generate internal representations of spaces and conduct reasoning based on these. Given a video input and a question, Yang et al[VSI-Bench] first construct a cognitive map consisting of object center positions in a grid format. Then the predicted map is leveraged to answer the question, achieving robust results regarding local distance awareness. Similarly, when exploring scenes in embodied tasks, a powerful internal representation is also useful to maintain past and current observations. For example, some works [ren2024explore, hu20253dllm] adopt 3D voxels as internal scene representation, on which the occupancy and exploration status are updated first. Afterwards, voxels are projected onto a 2D semantic map [ren2024explore]. For the unexplored areas, this map is utilized to predict probabilities over three candidate directions and to estimate whether further exploration is worthwhile based on current observations. (3) Attention re-distribution. During reasoning, there exists an imbalance between visual and textual attention. Specifically, image tokens constitute 90% of the input sequence but only account for 10% of the total attention, leading to misalignment between the actual spatial layout of objects in images and instruction prior in texts [chen2025spatial]. To deal with this imbalance, AdaptVis [chen2025spatial] introduces a dynamic attention redistribution strategy that leverages output logits as guidance. When the guidance is low, the attention distribution is smoothed, encouraging the model to explore a broader range of inputs. In contrast, when the guidance signal is high, the attention distribution is sharpened, directing the model to focus on key objects.
@@
External enhancement. External enhancement supplements VLMs with knowledge beyond their pretrained parameters to understand illogical or uncommon spatial relationships. One representative direction is the use of multi-agent system. Marsili et al. [marsili2025visual] designs a cooperative system where different agents interact in a code policy manner: API generation agents first decompose the query into sub-questions, while program synthesis agents receive these sub-questions and generate code to solve each problem. Another line of work is Retrieval Augmented Generation (RAG). For example, Yu et al[yu2024rag] retrieves templates for subjects-objects relationships as well as other spatial descriptions, and then integrates these templates into the inference stage via in-context learning.
@@ 4.1.1 Attributes
Attributes of objects refer to characteristics that remain relatively stable and are unlikely to change in the short term. These typically include shape, color, size, and quantity. The shape of an object is essential for distinguishing between different objects and understanding scenes. Object shape recognition tasks typically begin with simple 2D geometric figures, which include two types of task setups: whole-shape recognition [GeoMeter, Shapeworld] and sub-shape recognition [VisOnlyQA]. MM-Bench [MM-Bench] increases the difficulty by requiring models to recognize object shapes in real images. Additionally, some datasets use multi-view images as input, such as MMSI-Bench [MMSI-Bench], which asks the model to deduce an object’s shape from multiple views, synthesizing the information to infer the shape from a fixed viewpoint. Color also plays a critical role in object recognition, and color-related tasks are generally simpler, requiring the model to identify the color of a specific object [MM-Bench, Spatial457, OpenEQA]. In contrast, perceiving the size of an object is more complex. Some studies present two objects for comparison, asking the model to determine which is taller, shorter, wider, narrower, larger, or smaller. Other research requires the model to estimate an object’s size quantitatively, which involves using commonsense knowledge for direct measurement [VSI-Bench, SpatialRGPT, Open3DVQA]. In OMNI3D-Bench [Omni3D-Bench], a reference object’s physical measurement is provided as a standard. Object quantity recognition [PhysBench, GeoMeter, Spatial457, SPHERE, SITE, OST-Bench, OSR-Bench, SpatialBench] is one of the more challenging tasks in static attribute understanding. It depends on distinguishing between objects based on attributes such as shape, color, and size. This makes hallucination phenomena more pronounced in tasks involving object quantity recognition.
@@ 4.1.2 States
In contrast to attributes, states refer to characteristics of objects that are subject to change. Common state attributes include an object’s posture, orientation, and its open/closed status. In OpenEQA [OpenEQA], the task requires the model to estimate the open/closed state of a microwave door, while other studies focus on determining an object’s front-facing orientation in space [Spatial-MM, Spatial457, Super-CLEVR-3D]. Additionally, some research [LEGO-Puzzles, OST-Bench] introduces a temporal dimension, requiring the model to understand how states evolve over time. For instance, this might involve describing changes in an object’s rotational angle from the past to the present, or transformations in its shape resulting from specific actions [EOC-Bench]. While the perception of state attributes typically involves describing the properties of a single object or category of objects, and thus falls under basic perception, it is more complex than understanding static attributes. This complexity arises from the diversity of state attributes and the need for commonsense reasoning.
@@ 4.2.1 Static Understanding
Spatial Relationships. Early works in the paper VSR[VSR] identified 66 types of spatial relationships e.g., under, in front of, facing. The test data consists of image-text pairs, where the text includes a description and a label indicating correctness, such as: <Caption: The cow is ahead of the person. Label: False. >. This requires VLMs to first localize the two objects in the image and then determine whether the abstract relation ‘ahead of’ holds. Later studies introduced additional positional relationships, such as occlusion, reach, and containment. The difficulty in these tasks often lies not in object recognition but in understanding specific spatial relationships. The model must establish connections between the spatial structural relations between objects and the corresponding relation terms.
@@
Object Localization. The object localization task requires the model to accurately identify the position of a specified target within an image. The specific localization methods include detecting the precise spatial coordinates of the target object, determining which bounding box best fits the target, or estimating the coordinates of key points on the target. The BLINK [BLINK] test sample provides an image along with a labeled bounding box for the target object, where the model chooses the bounding box that best aligns with the target from multiple options. The SITE [SITE] involves the model locating a specific object based on a description or hint and estimating the coordinates of its 2D bounding box. Furthermore, the STI-Bench [STI-Bench] requires the model to localize the 3D bounding box of the target object within a video. Object localization tasks are also included in other benchmarks [OpenEQA, MM-Bench, SpatialBench, Visual-Cot, OpenEQA]. The challenge in these tasks lies in the model’s need to possess strong object detection capabilities while also handling various complexities in the background and interference, such as partial occlusion and similar objects.
@@
Distance Measuring. Distance measuring tasks aim to estimate the distance between objects in an image or the distance from an object to the camera, which is also referred to as depth estimation. Distance measuring typically involves two forms: qualitative judgment of proximity and quantitative numerical estimation. For instance, in the paper BLINK [BLINK], the task requires the model to determine which of two specified points is closer to the camera; in SPAR-Bench [SPAR-Bench], the task asks the model to judge which object is closer to or farther from a reference point (such as another object or an observer) based on the 3D center point position of the objects. In addition to qualitative distance analysis, SpatialVLM [SpatialVLM] employs an automated data construction pipeline to obtain distance data, enabling quantitative distance estimation. In the paper SpatialBench [SpatialBench], desktop-level distance estimation is introduced for use in embodied scenarios. Subsequently, many studies have incorporated distance measurement as a crucial task in spatial reasoning [Q-Spatial, SpatialRGPT, All-Angles-Bench, SPACE, VSTiBench, STI-Bench, VSI-Bench, RoomSpace, SPHERE, Proximity-110K, Omni3D-Bench, Open3DVQA, VSI-100K, SpatialScore, PhysBench, SITE]. The challenge of this task lies in the need for the model to infer spatial relationships using implicit cues such as structural information in the scene, object size ratios, and physical knowledge. Furthermore, varying camera parameters may introduce misguidance in distance perception.
@@
Spatial Compatibility Reasoning. Spatial compatibility reasoning tasks require VLMs to determine whether one object can fit the spatial properties or another after thoroughly understanding the object’s shape, size, and spatial relationships. In real-world scenarios, this includes determining whether an object can fit into a given space [OpenEQA] (’Can another cookie jar fit on the cookie jar shelf?’). This task assesses the model’s reasoning abilities concerning spatial constraints and physical adaptation relationships. Additionally, OmniSpatial [OmniSpatial] introduces reasoning about motion processes, such as determining whether an object can pass through a narrow passage. Some synthetic datasets also require the model to understand layout constraints and graphical combination rules in a 2D space, enabling it to provide fitting solutions. The difficulty of this task lies in the model’s need to not only perceive the size and shape of objects but also reason about physical world common sense, such as “an object too large cannot fit into a small space” or “some uniquely shaped objects cannot be stacked together”.
@@
Trajectory Description. The goal of the trajectory description task is to identify and describe the dynamic changes in the object or camera pose over a time sequence based on visual inputs, generating structured language descriptions. In this context, the trajectory refers to the continuous recording of the camera’s pose over time. The descriptions can be divided into two categories based on the observer (camera viewpoint) and the observed world (objective world): one describes the camera’s orientation and pose, and the other describes the objects within the observed world. In STI-Bench [STI-Bench], the model is required to infer the position and orientation of the camera at a specific time based on video input and initial pose information. In the Ego-ST Bench [Ego-ST], the test requires the model to qualitatively describe changes in direction, such as “turn left first and then turn right”. Additionally, multi-view reasoning tasks [MMSI-Bench, OmniSpatial, SpatialScore, BLINK, STI-Bench, VSTiBench, CameraBench, All-Angles-Bench, SPARE3D] involve numerous estimations of quantified camera parameters. Another category of trajectory description pertains to the observed world, involving tasks like recognizing changes in the direction of other objects in the video, such as “the car changes from moving straight to turning left”. In the VSI-Bench [VSI-Bench], a task is designed that requires the model to determine the order of appearance of objects in a spatiotemporal sequence. Similarly, in the SPACE [SPACE], multiple landmarks are introduced, and the model is required to determine their sequence. The challenges in these tasks arise from two key issues: first, the model must infer 3D spatial from 2D RGB inputs, which involves geometric principles such as perspective projection, a problem that cannot be resolved by simple image feature matching alone. Second, the estimation errors accumulate as the video sequence length increases, requiring the model to have dynamic error correction capabilities. This is a significant challenge in video pose estimation.
@@
Velocity and Acceleration Estimation. In dynamic spatial reasoning tasks, models are required to estimate motion parameters, such as velocity and acceleration, which are crucial for subsequent planning and decision-making. The OmniSpatial [OmniSpatial] requires models to qualitatively assess the speed and acceleration of objects, like “Which car, A or B, is moving faster?” or “Is the acceleration of the ball on the slope increasing?”. In contrast, STI-Bench directly asks models to estimate the numerical values of velocity and acceleration from video, such as “What is the average speed of the camera?” or “How quickly is the ball accelerating?”. Additionally, STI-Bench [STI-Bench] categorizes tests into three different environments: desktop, indoor, and outdoor, based on varying accuracy requirements for motion parameter estimations. In DynSuperCLEVR [DynSuperCLEVR], the test introduces two additional settings: future prediction and counterfactual reasoning. Future prediction requires the model to infer the future motion state of an object based on its current speed and acceleration, such as “How fast will the blue sphere be moving after 2 seconds?”. Counterfactual reasoning, on the other hand, asks the model to reason about hypothetical scenarios where speed or acceleration changes, such as “If the green cylinder’s acceleration were doubled, when would it collide with the wall? The task of estimating velocity and acceleration requires the model to integrate displacement changes and temporal information within a 3D space while relying on an understanding of physical laws. The challenges of this task arise from two key issues: first, velocity and acceleration are multi-stage dynamic attributes. Estimating speed involves calculating displacement, followed by the estimation of acceleration, which creates a longer inference chain. Second, counterfactual reasoning and future prediction tasks require the model to reconstruct scenarios, involving complex “what-if” logic chains that are more difficult than simple factual queries.
@@
Spatial Imagination. In this paper, we define spatial imagination tasks as those require reasoning from a hypothetical viewpoint different from the input perspective, given a visual input. It is important to note that some works [OmniSpatial, Open3DVQA, ViewSpatial-Bench] refer to the camera viewpoint as the “egocentric perspective”, while others [SPHERE, COMFORT] refer to it as the “allocentric perspective.” To avoid ambiguity, we use the term “hypothetic perspective” to distinguish it from the camera viewpoint, which is the perspective directly provided to the model. The inputs can be a single image, requiring the model to reason from a hypothetical perspective. For example, “If the waiter is on the side of the vase, then who is on his left?” [OmniSpatial] The SQA3D requires the model to first locate a hypothetical viewpoint in the video based on the description and then perform a question-answering task, such as “Sitting at the edge of the bed and facing the sofa, can I go straight to the coffee table in front of me?” The MINDCUBE [MINDCUBE] considers reasoning from limited views, aiming to assess the model’s ability to mentally model space, such as “If you move from view 1 to view 2, what is the furthest from you?” There is a significant body of research [SITE, VSI-Bench, Spatial-MM, SPAR-Bench] in this area, some of which also accounts for spatial understanding differences across various viewpoints and language conventions [COMFORT], or conducts reasoning from synthetic collections of images [SRBench]. The naming of spatial imagination tasks varies across studies, SQA3D [SQA3D] labeling it as situational question-answering, and SRBench [SRBench] referring to it as mental rotation. Spatial imagination tasks are not simply direct perception of static object attributes but require the model to construct spatial representations of objects in the mind and perform viewpoint transformations, involving dynamic operations on spatial structures. Thus, the challenge of spatial imagination lies in the need of the model to overcome the limitations of the input perspective, build accurate topological and spatial structural relationships from limited viewpoint inputs and imagine visual cues from the hypothetical perspective.